The value of the rice grows according to the function, 𝑉=𝑘𝑒2√𝑡, how long should the dealer store the rice.


1
Expert's answer
2021-05-04T11:15:54-0400

Given

"V=Ke^{2\sqrt t}"

Rice can be stored in a way it maximized the present value of rice, if the interest rate is r.

 Then, the present value is given as:

"P=V.e^{-rt}"

"P=Ke^{2\sqrt {t}} \times e^{-rt}"

Apply log on both sides as follows: 

"lnP=lnk+2\sqrt t-rt"


"\frac{\delta in P}{\delta t}=\frac{2}{2\sqrt t}-r=0"


"\frac{1}{\sqrt t}=r"


"\frac{1}{r}=\sqrt t"


Optimal time "t^*=\frac{1}{r^2}"


In the second order derivative

"\frac{\delta ^2 v}{\delta ^2 t}=v\frac{\delta}{\delta t}(t^{\frac{-1}{2}}-r)\frac{\delta v}{\delta t}"


Here, the product rule is used;

Then, the second term vanishes when: "\frac{\delta v}{\delta t}=0"

Thus,

"\frac{\delta v}{\delta t}=\frac{-v}{2}\sqrt{t^3}<0"

This satisfies the second order condition for maximum.

Thus the dealer should store the rice up to "t=\frac{1}{r^2}"